HIGH RESOLUTION LASER SPECTROSCOPY OF THE TITANIUM MONOHALIDES TiCl AND TiBr.

\underline{A. G. ADAM} and SCOTT HOPKINS, Department of Chemistry, University of New Brunswick, Fredericton, NB, Canada E3B 6E2.

High resolution laser spectra of TiCl and TiBr have been taken in the blue region of the spectrum. The titanum monohalide molecules were produced by laser vaporization of a Ti rod followed by reaction with either CHCl_{3} or CH_{3}Br in a pulsed supersonic jet source. The TiCl molecule was recently studied by FT spectroscopy^{a} where the electronic transition was identified as ^{4}Γ - ^{4}Φ consistent with that of TiF. In the paper, band-by-band case (c) rotational analyses of the (0,0) ^{4}Γ_{5/2} - ^{4}Φ_{3/2}, ^{4}Γ_{7/2} - ^{4}Φ_{5/2} and ^{4}Γ_{9/2} - ^{4}Φ_{7/2} subbands were given. The authors observed no subbands which would provide the information crucial to determining the spin-orbit spacing. We have observed this same TiCl electronic transition with our apparatus and report here the observation of the (0,0) ^{4}Γ_{5/2} - ^{4}Φ_{5/2} and ^{4}Γ_{7/2} - ^{4}Φ_{7/2} subbands, in addition to those given above, and provide a case (a) molecular analysis. The same electronic transition for TiBr has also been recorded. Although the spectrum is much more congested than that of TiCl, we have managed to assign a number of the subbands and are also able to provide a case (a) analysis for this molecule.

aT. Imajo, D. Wang, K. Tanaka, and T. Tanaka, J. Mol. Spectrosc. \underline{203}, 216 (2000).